OpenStudy (anonymous):

A person running an 800 meter race averages 130 meters per minutes for the first 3/4 of the race. the average speed for the remainder of the race is 145 meters per minute. what is the person's average speed for entire 800 meters rounded to the nearest whole number?

OpenStudy (anonymous):

Set up two equations using x=x0+v*t 1. 600=0+130*t => t=4.61538 2.800=600+145*t => t=1.37931 t total = 5.99 or 6 min 3. 800=0+Vavg*6 => Vavg = 133.33 meter/min